Compute the n-th differential coefficient of \\(y=x\\log_{e}x\\)\n
1
Expert's answer
2019-10-25T10:31:19-0400

"y=(x\ln x)^n" ;

"y'=n(\ln x+1)(x\ln x)^{n-1}" ;

"y''=n(n-1)(\ln x+1)(x\ln x)^{n-2}+n\frac{1}{x}(x\ln x)^{n-1}=n(x\ln x)^{n-1}\left(\frac{(n-1)(\ln x+1)}{x\ln x}+\frac{1}{x}\right)" ;


"y^{(n)}=n!\ln^nx"